OpenStudy (anonymous):

I just need explanation: Question: In the interval 0

OpenStudy (anonymous):

It says on the book that: \[v=L \frac{ di }{ dt } =15.0\cos50t (V)\] \[p=vi=75.0\sin 100t (W)\] \[w _{l}=\int\limits_{t}^{0}pdt=0.75(1-\cos100t)(J)\]

OpenStudy (anonymous):

\[v = L \frac{ di }{ dt } = 30.10^{-3} . \frac{ d(10\sin50t) }{ dt } = 30.10^{-3} .10.50\cos(50t)=15\cos(50t)\]

OpenStudy (anonymous):

\[p = vi = 15\cos50t . 10\sin50t = 150.0.5.\sin100t = 75\sin100t (W)\]

OpenStudy (anonymous):

so sin50*cos50 is sine 100?

OpenStudy (anonymous):

\[w_t = \int\limits_{0}^{t}pdt = \int\limits_{0}^{t}75\sin(100t)dt = \frac{ 75.(-\cos100t) }{ 100 } - \frac{ 75.(-\cos0) }{ 100 } = 0.75(1-\cos100t)\] (J) Hint: sin2x = 2sinx.cosx. One note that, the formula for energy should integrate from 0 --> t, not t --> 0 as your book said

OpenStudy (anonymous):

your post is cut

OpenStudy (anonymous):

Yes, you can solve the rest by yourself, it cut due to limit of page :D, maybe a good way to try to learn from the rest

OpenStudy (anonymous):

yes,,thanks

OpenStudy (anonymous):

btw..it is only sin50*cos50 not 2sin50*cos50

OpenStudy (anonymous):

No, you can see my answer, I put 0.5 factor, which means, sin50t*cos50t = 0.5sin100t, which means similar to sin100t = 2sin50t*cos50t

OpenStudy (anonymous):

without 0.5, you can not get the correct answer ;) Cheers

OpenStudy (anonymous):

ohh....ok thank you